jane works due north of home. Her husband Juan works due east. They leave for work at the same time.by the time jane is 3 miles from home, the distance between them is one more than juan's distance from home. How far from home is juan?
Answer by mananth(16949)

Since Jane is due north and Juan is due east they are moving at right angles to each other.
So Pythagoras theorem has to be applied.
at a given time jane is 3 miles from home
Let Juan be x miles from home
they are at a distance of (x+1) miles
the two legs are 3 & x
The hypotenuse is (x+1)
x^2+3^2=(x+!)^2
x^2+9= x^2+2x+1
x^2 cancels off
9=2x+1
2x=9-1
2x=8
/2
x= 4
Juan is 4 miles from home